OpenStudy (anonymous):

Find the first 3 terms, in ascending powers of x, of the binomial expansion of (2+x)^6 , giving each term in the simplest form?

OpenStudy (anonymous):

1st term: 2^6 = 64 2nd term: 6*2^5*x = 192x 3rd term: 6*5*2^4*x^2 = 480x^2

OpenStudy (anonymous):

(3x+6x+2x) = -4+5+8 11x = 9 x = 9/11